OpenStudy (anonymous):

HURRY!!!! Which of the following are identities? Check all that apply. a. sin2x = 1 + cos2x b. sin2x + cos2x = 1 c. tan2x = 1 - sec2x d. cot2x + 1 = csc2x

OpenStudy (anonymous):

I think B works, but what else?

OpenStudy (ash2326):

\[\sin^2x+\cos^2 x=1\] \[\cot^2x+1=cosec^2 x\]